The following monthly data are available for Crane Company which produces only one product: Selling price per unit, $44; Unit variable expenses, $14; Total fixed expenses, $42000; Actual sales for the month of June, 7000 units. How much is the margin of safety for the company for June?
Answer #1

Break even unit = Fixed cost/Contribution margin per unit = 42000/(44-14) = 1400 Units

Margin of safety = 7000-1400 = 5600 Units

Margin of safety = (7000-1400)*44 = $246400